Anna's age is twice that of David's and half that of Preeya's. The total of all their ages is 21. How old is Preeya?

Answer by TimothyLamb(4379)

x = anna age
y = david age
z = preeya age
---
x + y + z = 21
x = 2y
z = 2x
---
put the system of linear equations into standard form
---
x + y + z = 21
x - 2y = 0
2x - z = 0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = anna age = 6
y = david age = 3
z = preeya age = 12
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---